# LongAudio-RAG: Event‑Grounded Question Answering over Multi‑Hour Long Audio

### v2

@@ Appendix H Dialogue evaluation using humans and GPT-4o as judge
We evaluated a dialogue system on 113 single‑turn QA pairs across three task categories detection, counting, and summarization using both a human evaluator and a GPT-4o based evaluator Achiam et al. (2023). For every response, each judge assigned 1–5 scores for language quality, correctness, reasoning coherence, and user experience; we averaged these four scores to obtain a per‑item composite and then averaged those composites within each category to produce the final category scores. Human evaluation yielded category means of 4.36, 4.25, and 3.95 (detection, counting, summarization), while GPT-4o based evaluation yielded 4.46, 3.67, and 3.47 for the same categories. The Pearson correlation between human and GPT per‑item composites across all 114 examples was 0.57, indicating moderate alignment. Despite comparatively lower outcomes in counting and summarization, performance is likely to improve with model scaling or targeted fine‑tuning within the size–accuracy budget. Overall, the system delivers strong conversational performance and coherent execution across categories
Appendix I Anomaly Detection Extension
Our framework naturally supports anomaly detection by treating anomalies as structured event-level signals that can be retrieved, analyzed, and explained using the same time-resolution and intent-routing components. When users ask anomaly-related questions, the system first classifies the query subtype (e.g., loudness anomaly, pitch anomaly, or abnormal start-time), retrieves the relevant events, and computes anomalies using a plug-in detector. In our implementation, we evaluate two forms of attribute-based anomalies such as loudness deviations from the normal range and temporal anomalies, where event start times fall outside expected intervals derived from historical clusters. The system produces a structured anomaly table and, when needed, incorporates guidance from user-provided instruction manuals before prompting the LLM to generate a grounded explanation. This allows the method to flexibly support any anomaly detection backend while maintaining faithful, event-grounded audio question answer capabilities.
We have implemented anomaly detection based on the loudness attribute and also irregular sound patterns. Figures 10 and 8 show an example with the flowchart analysis pipeline that can be integrated into our proposed system. Figure 9 shows how we can integrate user manuals and other custom logs into our system for more personalized responses to target the specific use-case.
Figure 7: Loudness anomaly detection
Figure 8: Frequency based anomaly detection
Figure 9: Anomaly detection processing for long audio QA
Appendix J Limitations
@@ Appendix I Anomaly Detection Extension
Appendix K Software Implementation
@@
Figure 10: Software Implementation Architecture
Appendix L User Interface for LongAudio-RAG
@@
Figure 11: LongAudio-RAG user interface
Appendix M Prompts used for intent based response generation
The prompts used for the response generation are shown in Figures 12 and 13.
Figure 12: Prompt for Summary category
Figure 13: Prompt for detection/counting based questions

### v3

@@ Appendix D Dialogue evaluation using humans and GPT-4o as judge
We evaluated a dialogue system on 113 single‑turn QA pairs across three task categories detection, counting, and summarization using both a human evaluator and a GPT-4o based evaluator Achiam et al. (2023). For every response, each judge assigned 1–5 scores for language quality, correctness, reasoning coherence, and user experience; we averaged these four scores to obtain a per‑item composite and then averaged those composites within each category to produce the final category scores. Human evaluation yielded category means of 4.36, 4.25, and 3.95 (detection, counting, summarization), while GPT-4o based evaluation yielded 4.46, 3.67, and 3.47 for the same categories. The Pearson correlation between human and GPT per‑item composites across all 113 examples was 0.57, indicating moderate alignment. Despite comparatively lower outcomes in counting and summarization, performance is likely to improve with model scaling or targeted fine‑tuning within the size–accuracy budget. An evaluator with validated industrial experience in audio editing and structured audio data collection oversaw all human assessments.
Appendix E Limitations
@@
Appendix F Software Implementation
@@
Appendix G User Interface for LongAudio-RAG